OpenStudy (anonymous):

graph the curves y=IxI-1, y=1-x^2. Identify the points of intersection and hence the area enclosed by the 2 curves.

myininaya (myininaya):

|x|-1=1-x^2 set equations = to find where they intersect |x|-1-1+x^2=0 |x|-2+x^2=0 x^2-|x|-2=0 if x>0, |x|=x if x<0, |x|=-x so lets assume the fist x>0 x^2-x-2=0 (x-2)(x+1)=0 x=2,x=-1 now lets assume the second x<0 x^2-(-x)-2=0 x^2+x-2=0 (x+2)(x-1)=0 x=-2, x=1

myininaya (myininaya):

so we have four points on intersection

myininaya (myininaya):

maybe i made a mistake

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

when you say consider x<0 , you should take the negative x soln

OpenStudy (anonymous):

so x=2 and x=-2

OpenStudy (anonymous):

are the ones u want

OpenStudy (anonymous):

wait, they arent even correct :|

myininaya (myininaya):

ok we can check all of our possible points of intersection we have the equation |x|-1=1-x^2 lets check x=1 |1|-1=1-(1)^2 0=0 so x=1 works

myininaya (myininaya):

x=-1 |-1|-1=1-(-1)^2 0=0 so -1 works x=-2 |-2|-1=1-(-2)^2 1=-3 -2 doesn't work |2|-1=1-2^2 1=-1 -2 doesn't work

myininaya (myininaya):

ok so after you get your possible points of intersection plug them back into the both equations to see if you get the same thing if you don't then they don't intersect there

myininaya (myininaya):

so they intersect at -1 and 1 according to what i did above

OpenStudy (anonymous):

now to find the area you have to find out which of the functions are greater at the intersection points

myininaya (myininaya):

now to find area \[\int\limits_{-1}^{1}(topfunction-bottomfunction)dx=\int\limits_{-1}^{1}((1-x^2)-(|x|-1) )dx\]

myininaya (myininaya):

to integrate \[\int\limits_{-1}^{1}|x| dx=\int\limits_{-1}^{0}-x dx+\int\limits_{0}^{1}x dx\]

myininaya (myininaya):

just in case you didn't know about that part

myininaya (myininaya):

do you need more help? do you understand? are there any questions? what are more ways to ask this question? lol

OpenStudy (anonymous):

yes,need help

OpenStudy (anonymous):

A=1 if my math is correct

myininaya (myininaya):

on what part

OpenStudy (anonymous):

after integrating...

OpenStudy (anonymous):

sub the values to get the area?

OpenStudy (anonymous):

i need guidance plz

myininaya (myininaya):

\[\int\limits_{-1}^{1}|x|dx=\frac{-x^2}{2} ,x=-1..0 +\frac{x^2}{2}, x=0..1=\frac{-(0)^2}{2}-\frac{-(-1)^2}{2}+\frac{1^2}{2}-\frac{0^2}{2}\] \[=\frac{1}{2}+\frac{1}{2}\] \[=1\] okay now lets look at the whole thing \[\int\limits_{-1}^{1}((1-x^2)-(|x|-1))dx=\int\limits_{-1}^{1}(2-x^2-|x|) dx=\int\limits_{-1}^{1}2dx-\int\limits_{-1}^{1}x^2dx-\int\limits_{-1}^{1}|x|dx\]

myininaya (myininaya):

\[(2x, x=-1..1)-(\frac{x^3}{3}, x=-1..1)-1\]

myininaya (myininaya):

\[[2(1)-2(-1)]-[\frac{1^3}{3}-\frac{(-1)^3}{3}]-1\]

myininaya (myininaya):

\[2+2-\frac{1}{3}+\frac{-1}{3}-1=4-\frac{2}{3}-1=3-\frac{2}{3}=\frac{3*3}{3}-\frac{2}{3}=\frac{9-2}{3}=\frac{7}{3}\] if i didn't make a mistake

myininaya (myininaya):

dang i made a mistake lol

myininaya (myininaya):

oh what did i let me look again

myininaya (myininaya):

looks good :)

myininaya (myininaya):

any questions? :)

OpenStudy (anonymous):

oh that's the final answer

myininaya (myininaya):

yes 7/3 is the final correct answer

OpenStudy (anonymous):

brilliant,thanks a lot.God bless